Let X, Y E [0, 1] be distributed according to the joint distribution Íxy (z, y) 6xy2 . Let -XY-3 . Find P(Z < 1 /2)
Answer #1

P(Z < 1/2)

= P(XY^3 < 1/2)

= 1 - P(Xy^3 > 1/2)

Now

P(Xy^3 > 1/2) =

= 0.53 0.5

2 0.75

2..2 0.75 ) dy = 0.25

required probability = 1- 0.25 = 0.75